OpenStudy (anonymous):

MEDAL & FAN ANYBODY ? If the height and radius of a cone are tripled, then the volume will be 9 times the original size. True False

OpenStudy (campbell_st):

a very simply solution to this is find the volume of 2 cones. 1. a cone radius 1 height 1 \[V = \frac{1}{3} \pi \times 1^2 \times 1 = \frac{1}{3} \pi\] then tripling these measures you will have 2. radius 3 and height 3 now find the volume of the 2nd cone anc compare them

OpenStudy (campbell_st):

or you can wait for someone to post an answer

OpenStudy (mrnood):

the volume of any shape is proportional to the CUBE of its dimensions so for instance if the dimension (eg radius) then is 1 a shape with radius 2 will have volume 2^3 times the original. The same applies for radius 3.....

OpenStudy (anonymous):

@campbell_st so i would multiply ?

OpenStudy (campbell_st):

what did you get for the volume of the cone where the measurements were tripled r = 3 and h = 3 \[V = \frac{1}{3} \times \pi \times 3^2 \times 3\]

OpenStudy (campbell_st):

and if you look carefully at the formula